OpenStudy (anonymous):

Help! Please! Find all solutions to the equation. sin^2 x + sin x = 0

OpenStudy (anonymous):

Pull out the common thingy like this 4 + 2 = 0 2(2+1) = 0

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

sinx(sinx + 1) = 0

OpenStudy (anonymous):

chaguanas's example showed squared and and un-squared values 4 + 2 = is 2^2 + 2 = so factoring out the common "thingy" 2 (2+1) = sin^2x and sin x so factor out sinx, like christiangeek did for you.

OpenStudy (anonymous):

When two things multiplied together equal zero, the only way for that to be possible is if one (or both) of them are zero. So sinx(sins + 1) will only be zero when sinx is zero or sins +1 is zero.

OpenStudy (anonymous):

(Sorry, my stupid computer keeps autocorrecting sinx to sins!

OpenStudy (anonymous):

You handle is "christian" geek.

OpenStudy (anonymous):

Ironic, isn't it?!

OpenStudy (anonymous):

yeah, so what do I do in order to solve this problem?

OpenStudy (anonymous):

you first factor out the common "thingy" sin x, giving you: sin x (sin x + 1) = 0 Then as Christiangeek explained. one of the factors has to be zero so that they equal zero when multiplied. One or both of the following has to be true: sin x = 0 sin x + 1 = 0 Solve each for x and those are your solutions.

OpenStudy (anonymous):

oh, i get it thanks!